OpenStudy (anonymous):

simplfy 5(3x+4)-11 5(3x+4)-11 =

OpenStudy (anonymous):

Distribute 5 to brackets first..

OpenStudy (anonymous):

so put 5 where the x is?

OpenStudy (anonymous):

x=-3/5 :)

OpenStudy (anonymous):

make shore to right the :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[15x +20 -11=0\] \[15x + 9 = 0\] \[15x = -9\] \[x =- \frac{ 9 }{ 15 }\] x = -3/5

OpenStudy (anonymous):

Don't give just answers @85295james .. It is against the Code Of Conduct of this site.. Please read this : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

OK sorry :(

OpenStudy (anonymous):

Distribute 5 to the brackets means Multiply 5 to the terms inside the brackets..

OpenStudy (anonymous):

i got 15x +9

OpenStudy (anonymous):

Like : For example we have 3(4x + 10) So multiply 3 with 4x and 10. 3(4x + 10) = 3*4x + 3*10 = 12x + 30 Like this multiply 5 with the terms inside that brackets.. Can you do that @ash0629 ??

OpenStudy (anonymous):

It is okay @85295james No need to say sorry.. Be careful next time.. Just back your answers up with your solutions too..

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

ok